# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ A.1.5 Implications for KSDs on Non-Euclidean Spaces
The core argument underpinning the proofs in Theorem 1 is that the function $\mathbf{z}\mapsto u_{p}(\mathbf{z},\mathbf{z})$ is unbounded on the support of the model, which in our case is $\mathbb{R}^{d}$. The proof per se does not rely on the support being an Euclidean space. Consequently, this argument can potentially be extended to KSDs constructed for models defined on other spaces, such as models for Riemannian data (Xu and Matsuda,, 2021), sequence data with varying dimensionality (Baum et al.,, 2023), functional data (Wynne et al.,, 2022), or censored time-to-event data (Fernandez et al.,, 2020).
A.2 Proof of Lemma 2
@@
$u_{p}(\mathbf{x},\mathbf{x})\;$ $=\;\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}^{2}h(0)+2\langle\mathbf{s}_{p,w}(\mathbf{x}),\;\nabla w(\mathbf{x})\rangle h(0)+\|\nabla w(\mathbf{x})\|_{2}^{2}h(0)-w(\mathbf{x})^{2}\nabla^{\top}\nabla h(0)$ (28)
@@
A.3 Proof of Theorem 3 and Related Preliminary Results
Overview of proof. We will show a general result in Proposition 14, which states that the robust-KSD test that rejects $H^{\mathrm{C}}_{0}:\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\Delta_{\theta}(\mathbb{X}_{n})\coloneqq\max(0,D(\mathbb{X}_{n})-\theta)>q_{B,1-\alpha}(\mathbb{X}_{n})$ is qualitatively robust for any $\theta\geq 0$. This immediately shows Theorem 3 by setting $\theta=0$. To show Proposition 14, we follow a similar approach in the proof of Theorem 1, where we will first show that the result holds conditionally on the number of contaminated data, and use a high-probability argument to complete the proof.
@@ A.3 Proof of Theorem 3 and Related Preliminary Results
  • Section A.3.1 shows Lemma 11, which states that the bootstrap quantiles $q_{B,1-\alpha}(\mathbb{X}_{n})$ and $q_{B,1-\alpha}(\mathbb{X}_{n}^{\ast})$ computed using any two samples $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ that differ by at most $m^{\prime}$ elements are close to each other with high probability.
  • Section A.3.2 shows Lemma 12, which says that the difference in the exceedance probabilities of $\Delta_{\theta}(\mathbb{X}_{n})$ and of $\Delta_{\theta}(\mathbb{X}_{n}^{\ast})$ is small.
  • Section A.3.3 shows Lemma 13, which says that the rejection probability of a robust-KSD test using $\mathbb{X}_{n}$ is close to that of a test using $\mathbb{X}_{n}^{\ast}$, assuming $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ differ by no more than $o(n^{1/2})$ elements. Its proof relies on Lemma 11 and Lemma 12.
  • Section A.3.4 uses Lemma 13 to show Proposition 14, which says that the robust-KSD test is qualitatively robust for any $\theta\geq 0$. This result immediately shows Theorem 3, the proof of which is also contained in this subsection.
@@
Assume $k$ is a tilted kernel satisfying the conditions in Lemma 2. Then there exists absolute constants $C_{1},C_{2}>0$ such that, for any $\delta>0$ and any (deterministic) realisations $\mathbb{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and $\mathbb{X}_{n}^{\ast}=\{\mathbf{x}_{i}^{\ast}\}_{i=1}^{n}$ that differ by at most $m^{\prime}$ elements, we have
[…]
@@
[…]
@@
Assume $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$ and that $k$ is a tilted kernel satisfying the conditions in Lemma 2. For any $\delta,\gamma>0$ and any integer $m^{\prime}\in[0,n]$, it holds that
@@
Assume $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$ and that $k$ is a tilted kernel satisfying the conditions in Lemma 2. For any integers $\theta\geq 0$, $\delta>0$, and any sequence $f_{n}=o(n^{1/2})$, there exists $n_{0}$ such that for any $n\geq n_{0}$, we have
@@
$\omega(m^{\prime})\;$ $\coloneqq\;\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\left(\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\right)-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\left(\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\right)\big{|}\;.$
@@
Assume $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$ and that $k$ is a tilted kernel satisfying the conditions in Lemma 2. If $\theta\geq 0$ and $\epsilon_{n}=o(n^{-1/2})$, then as $n\to\infty$,
[…]
@@ A.3.1 Proof of Lemma 11
$n^{2}\big{|}D^{2}_{\mathbf{W}}(\mathbb{X}_{n})-D^{2}_{\mathbf{W}}(\mathbb{X}_{n}^{\ast})\big{|}\;$ $=\;\big{|}\mathbf{V}_{2}^{\top}U\mathbf{V}_{2}+2\mathbf{V}_{1}^{\top}U\mathbf{V}_{2}-\mathbf{V}_{2}^{\top}U^{\prime}\mathbf{V}_{2}-2\mathbf{V}_{1}^{\top}U^{\prime}\mathbf{V}_{2}\big{|}$
@@
where the last line follows from Cauchy-Schwarz inequality applied to the $U$-weighted inner product $(\mathbf{x},\mathbf{x}^{\prime})\mapsto\mathbf{x}^{\top}U\mathbf{x}^{\prime}$ and the $U^{\prime}$-weighted inner product $(\mathbf{x},\mathbf{x}^{\prime})\mapsto\mathbf{x}^{\top}U^{\prime}\mathbf{x}^{\prime}$, which are well-defined since $U,U^{\prime}$ are positive semi-definite. The proof proceeds by bounding each term separately. We discuss how to bound $|\mathbf{V}_{1}U\mathbf{V}_{1}|$ and $|\mathbf{V}_{2}U\mathbf{V}_{2}|$, and the argument for $U^{\prime}$ is identical. We first define the $n\times n$ matrix with $(A_{1})_{ij}=u_{ij}$ for $i,j\leq m$ and $(A_{1})_{ij}=0$, so that we can write $\mathbf{V}_{1}^{\top}U\mathbf{V}_{1}=\mathbf{W}^{\top}A_{1}\mathbf{W}$. We can apply the Hanson-Wright lemma (Lemma 10) to conclude that there exists positive constants $C$ such that, for any $\delta>0$ and any almost-sure sequence $\mathbb{X}_{n}$, the following event occurs with probability at least $41-\delta/$ 4,
@@
where the second last inequality holds since $|u_{ij}|\leq\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}|u_{p}(\mathbf{x},\mathbf{x}^{\prime})|\leq\tau_{\infty}$, and the last line holds since $m\leq n$. Similarly, defining the matrix $A_{2}$ with $(A_{2})_{ij}=0$ for $i,j\leq m$ and $(A_{2})_{ij}=u_{ij}$ otherwise, we can write $\mathbf{V}_{2}^{\top}U\mathbf{V}_{2}=\mathbf{W}^{\top}A_{2}\mathbf{W}$, and the same argument as before shows that the following holds with probability at least $41-\delta/$ 4,
@@
Combining (32) and (33), we conclude that the following holds with probability at least $21-\delta/$ 2,
@@
The same argument shows that the last two terms of (31) can be bounded by the same quantity on an event with probability at least $21-\delta/$ 2. To conclude, we have shown that with probability at least $1-\delta$,
@@
$|q_{\infty,1-\alpha}(\mathbb{X}_{n})-q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})|\;\leq\;\max(q_{\infty,1-\alpha}(\mathbb{X}_{n}),\;q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast}))\;$ $\leq\;q_{\infty,1-\alpha}(\mathbb{X}_{n})+q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\;\leq\;\rho_{m,n,\delta}^{\frac{1}{2}}\;,$
@@ A.3.2 Proof of Lemma 12
where the last line holds since $\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\leq\tau_{\infty}$ under the assumed kernel conditions due to Lemma 2. We then bound $|S_{m}|$ and $|S_{m}^{\ast}|$ with high probability. When $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X}^{\ast})\|_{2}]<\infty$, we have $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ as argued in the paragraph before (11), so it is straightforward to see that $\mathbb{E}_{\mathbf{X}_{i}^{\ast}\sim P}[T_{i}]=0$ for all $i$, and thus $S_{m}$ is a sum of zero-mean i.i.d. random variables $T_{i}$. Therefore,
@@
where the last step holds again since $\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\leq\tau_{\infty}$. For any $\delta>0$, Chebyshev’s inequality implies that the event $\mathcal{A}_{1}\coloneqq\{|S_{m}|\leq\sqrt{m}(n-m)\tau_{\infty}/\sqrt{\delta/2}\}$ occurs with probability at least $21-\delta/$ 2. In other words, $|S_{m}|$ can be upper bounded on the high-probability event $\mathcal{A}_{1}$. A similar argument applied to $S_{m}^{\ast}$ shows that the event $\mathcal{A}_{2}\coloneqq\big{\{}|S_{m}^{\ast}|\leq\sqrt{m}(n-m)\tau_{\infty}/\sqrt{\delta/2}\}$ also occurs with probability at least $21-\delta/$ 2. On $\mathcal{A}\coloneqq\mathcal{A}_{1}\cap\mathcal{A}_{2}$, which occurs with probability at least $1-\delta$, we have by (35) that
@@
$\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P}\big{(}D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>\gamma\big{)}\;$ $=\;\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P}\big{(}\{D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>\gamma\}\cap\mathcal{A}\big{)}+\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P}\big{(}\{D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>\gamma\}\cap\mathcal{A}^{\mathsf{c}}\big{)}$

### v4

@@ A.1.5 Implications for KSDs on Non-Euclidean Spaces
The core argument underpinning the proofs in Theorem˜1 is that the function $\mathbf{z}\mapsto u_{p}(\mathbf{z},\mathbf{z})$ is unbounded on the support of the model, which in our case is $\mathbb{R}^{d}$. The proof per se does not rely on the support being an Euclidean space. Consequently, this argument can potentially be extended to KSDs constructed for models defined on other spaces, such as models for Riemannian data (Xu and Matsuda,, 2021), sequence data with varying dimensionality (Baum et al.,, 2023), functional data (Wynne et al.,, 2022), or censored time-to-event data (Fernandez et al.,, 2020).
A.2 Proof of Lemma˜2
@@
$u_{p}(\mathbf{x},\mathbf{x})$
$=\;\|\mathbf{s}_{p,w}(\mathbf{x})\|_{2}^{2}h(0)+2\langle\mathbf{s}_{p,w}(\mathbf{x}),\;\nabla w(\mathbf{x})\rangle h(0)+\|\nabla w(\mathbf{x})\|_{2}^{2}h(0)-w(\mathbf{x})^{2}\nabla^{\top}\nabla h(0)$ (28)
@@
A.3 Proof of Theorem˜3 and Related Preliminary Results
Overview of proof. We will show a general result in Proposition˜14, which states that the robust-KSD test that rejects $H^{\mathrm{C}}_{0}:\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ if $\Delta_{\theta}(\mathbb{X}_{n})\coloneqq\max(0,D(\mathbb{X}_{n})-\theta)>q_{B,1-\alpha}(\mathbb{X}_{n})$ is qualitatively robust for any $\theta\geq 0$. This immediately shows Theorem˜3 by setting $\theta=0$. To show Proposition˜14, we follow a similar approach in the proof of Theorem˜1, where we will first show that the result holds conditionally on the number of contaminated data, and use a high-probability argument to complete the proof.
@@ A.3 Proof of Theorem˜3 and Related Preliminary Results
  • Section˜A.3.1 shows Lemma˜11, which states that the bootstrap quantiles $q_{B,1-\alpha}(\mathbb{X}_{n})$ and $q_{B,1-\alpha}(\mathbb{X}_{n}^{\ast})$ computed using any two samples $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ that differ by at most $m^{\prime}$ elements are close to each other with high probability.
  • Section˜A.3.2 shows Lemma˜12, which says that the difference in the exceedance probabilities of $\Delta_{\theta}(\mathbb{X}_{n})$ and of $\Delta_{\theta}(\mathbb{X}_{n}^{\ast})$ is small.
  • Section˜A.3.3 shows Lemma˜13, which says that the rejection probability of a robust-KSD test using $\mathbb{X}_{n}$ is close to that of a test using $\mathbb{X}_{n}^{\ast}$, assuming $\mathbb{X}_{n}$ and $\mathbb{X}_{n}^{\ast}$ differ by no more than $o(n^{1/2})$ elements. Its proof relies on Lemma˜11 and Lemma˜12.
  • Section˜A.3.4 uses Lemma˜13 to show Proposition˜14, which says that the robust-KSD test is qualitatively robust for any $\theta\geq 0$. This result immediately shows Theorem˜3, the proof of which is also contained in this subsection.
@@
Assume $k$ is a tilted kernel satisfying the conditions in Lemma˜2. Then there exists absolute constants $C_{1},C_{2}>0$ such that, for any $\delta>0$ and any (deterministic) realisations $\mathbb{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and $\mathbb{X}_{n}^{\ast}=\{\mathbf{x}_{i}^{\ast}\}_{i=1}^{n}$ that differ by at most $m^{\prime}$ elements, we have
$\Pr\nolimits_{\mathbf{W}}\left(\left|D^{2}_{\mathbf{W}}(\mathbb{X}_{n})-D^{2}_{\mathbf{W}}(\mathbb{X}_{n}^{\ast})\right|\;\leq\;\frac{\tau_{\infty}\epsilon_{m,n}^{\frac{1}{2}}}{n}\left(C_{1}+C_{2}\log\left(\frac{8}{\delta}\right)\right)\right)\;\geq\;1-\delta\;,$
@@
$\Pr\nolimits_{\mathbf{W}}\left(\left|q_{\infty,1-\alpha}(\mathbb{X}_{n})-q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\right|\;\leq\;\frac{\tau_{\infty}^{\frac{1}{2}}\epsilon_{m,n}^{\frac{1}{4}}}{n^{\frac{1}{2}}}\sqrt{C_{1}+C_{2}\log\left(\frac{8}{\delta}\right)}\;\right)\geq\;1-\delta\;.$
@@
Assume $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$ and that $k$ is a tilted kernel satisfying the conditions in Lemma˜2. For any $\delta,\gamma>0$ and any integer $m^{\prime}\in[0,n]$, it holds that
@@
Assume $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$ and that $k$ is a tilted kernel satisfying the conditions in Lemma˜2. For any integers $\theta\geq 0$, $\delta>0$, and any sequence $f_{n}=o(n^{1/2})$, there exists $n_{0}$ such that for any $n\geq n_{0}$, we have
@@
$\omega(m^{\prime})\;$ $\coloneqq\;\big{|}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\left(\Delta_{\theta}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>q_{\infty,1-\alpha}(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})\right)$
$\;\quad\quad-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\left(\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\right)\big{|}\;.$
@@
Assume $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$ and that $k$ is a tilted kernel satisfying the conditions in Lemma˜2. If $\theta\geq 0$ and $\epsilon_{n}=o(n^{-1/2})$, then as $n\to\infty$,
$\sup_{Q\in\mathcal{P}(P;\epsilon_{n})}\big{|}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})>q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n}^{\ast})>q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}\;\to\;0\;,$
@@ A.3.1 Proof of Lemma˜11
$n^{2}\big{|}D^{2}_{\mathbf{W}}(\mathbb{X}_{n})-D^{2}_{\mathbf{W}}(\mathbb{X}_{n}^{\ast})\big{|}$
$=\;\big{|}\mathbf{V}_{2}^{\top}U\mathbf{V}_{2}+2\mathbf{V}_{1}^{\top}U\mathbf{V}_{2}-\mathbf{V}_{2}^{\top}U^{\prime}\mathbf{V}_{2}-2\mathbf{V}_{1}^{\top}U^{\prime}\mathbf{V}_{2}\big{|}$
@@
where the last line follows from Cauchy-Schwarz inequality applied to the $U$-weighted inner product $(\mathbf{x},\mathbf{x}^{\prime})\mapsto\mathbf{x}^{\top}U\mathbf{x}^{\prime}$ and the $U^{\prime}$-weighted inner product $(\mathbf{x},\mathbf{x}^{\prime})\mapsto\mathbf{x}^{\top}U^{\prime}\mathbf{x}^{\prime}$, which are well-defined since $U,U^{\prime}$ are positive semi-definite. The proof proceeds by bounding each term separately. We discuss how to bound $|\mathbf{V}_{1}U\mathbf{V}_{1}|$ and $|\mathbf{V}_{2}U\mathbf{V}_{2}|$, and the argument for $U^{\prime}$ is identical. We first define the $n\times n$ matrix with $(A_{1})_{ij}=u_{ij}$ for $i,j\leq m$ and $(A_{1})_{ij}=0$, so that we can write $\mathbf{V}_{1}^{\top}U\mathbf{V}_{1}=\mathbf{W}^{\top}A_{1}\mathbf{W}$. We can apply the Hanson-Wright lemma (Lemma˜10) to conclude that there exists positive constants $C$ such that, for any $\delta>0$ and any almost-sure sequence $\mathbb{X}_{n}$, the following event occurs with probability at least $1-\delta/4$,
@@
where the second last inequality holds since $|u_{ij}|\leq\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}|u_{p}(\mathbf{x},\mathbf{x}^{\prime})|\leq\tau_{\infty}$, and the last line holds since $m\leq n$. Similarly, defining the matrix $A_{2}$ with $(A_{2})_{ij}=0$ for $i,j\leq m$ and $(A_{2})_{ij}=u_{ij}$ otherwise, we can write $\mathbf{V}_{2}^{\top}U\mathbf{V}_{2}=\mathbf{W}^{\top}A_{2}\mathbf{W}$, and the same argument as before shows that the following holds with probability at least $1-\delta/4$,
@@
Combining (32) and (33), we conclude that the following holds with probability at least $1-\delta/2$,
@@
The same argument shows that the last two terms of (31) can be bounded by the same quantity on an event with probability at least $1-\delta/2$. To conclude, we have shown that with probability at least $1-\delta$,
@@
$|q_{\infty,1-\alpha}(\mathbb{X}_{n})-q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})|\;\leq\;\max(q_{\infty,1-\alpha}(\mathbb{X}_{n}),\;q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast}))\;$ $\leq\;q_{\infty,1-\alpha}(\mathbb{X}_{n})+q_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})$
$\leq\;\rho_{m,n,\delta}^{\frac{1}{2}}\;,$
@@ A.3.2 Proof of Lemma˜12
where the last line holds since $\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\leq\tau_{\infty}$ under the assumed kernel conditions due to Lemma˜2. We then bound $|S_{m}|$ and $|S_{m}^{\ast}|$ with high probability. When $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X}^{\ast})\|_{2}]<\infty$, we have $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ as argued in the paragraph before (11), so it is straightforward to see that $\mathbb{E}_{\mathbf{X}_{i}^{\ast}\sim P}[T_{i}]=0$ for all $i$, and thus $S_{m}$ is a sum of zero-mean i.i.d. random variables $T_{i}$. Therefore,
@@
where the last step holds again since $\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\leq\tau_{\infty}$. For any $\delta>0$, Chebyshev’s inequality implies that the event $\mathcal{A}_{1}\coloneqq\{|S_{m}|\leq\sqrt{m}(n-m)\tau_{\infty}/\sqrt{\delta/2}\}$ occurs with probability at least $1-\delta/2$. In other words, $|S_{m}|$ can be upper bounded on the high-probability event $\mathcal{A}_{1}$. A similar argument applied to $S_{m}^{\ast}$ shows that the event $\mathcal{A}_{2}\coloneqq\big{\{}|S_{m}^{\ast}|\leq\sqrt{m}(n-m)\tau_{\infty}/\sqrt{\delta/2}\}$ also occurs with probability at least $1-\delta/2$. On $\mathcal{A}\coloneqq\mathcal{A}_{1}\cap\mathcal{A}_{2}$, which occurs with probability at least $1-\delta$, we have by (35) that
@@
$\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P}\big{(}D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>\gamma\big{)}$
$=\;\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P}\big{(}\{D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>\gamma\}\cap\mathcal{A}\big{)}+\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P}\big{(}\{D(\mathbb{X}_{m}^{\ast}\cup\mathbb{Z}_{m^{\prime}})>\gamma\}\cap\mathcal{A}^{\mathsf{c}}\big{)}$